OpenStudy (anonymous):

(3t^7 − 8t^3 + 3) − (8t^7 − 3t^3 + 8) What is the answer to this, i tried but the computer kept saying it was wrong .

OpenStudy (bossvideogamer21):

what was the answer u put in

OpenStudy (bossvideogamer21):

it may be right but if it is u have to check ur spelling

OpenStudy (precal):

(3t^7 − 8t^3 + 3) − (8t^7 − 3t^3 + 8) 3t^7-8t^3+3-8t^7+3t^3-8 combine like terms

OpenStudy (johnn.):

Let's simplify step-by-step. 3t7−8t3+3−(8t7−3t3+8) Distribute the Negative Sign: =3t7−8t3+3+−1(8t7−3t3+8) =3t7+−8t3+3+−1(8t7)+−1(−3t3)+(−1)(8) =3t7+−8t3+3+−8t7+3t3+−8 Combine Like Terms: =3t7+−8t3+3+−8t7+3t3+−8 =(3t7+−8t7)+(−8t3+3t3)+(3+−8) =−5t7+−5t3+−5 Answer: =−5t7−5t3−5

OpenStudy (anonymous):

Dang, i had +5 at the end. thanks though